# ShiftAddLLM: Accelerating Pretrained LLMs via Post-Training Multiplication-Less Reparameterization

### v3

@@ Abstract
Large language models (LLMs) have shown impressive performance on language tasks but face challenges when deployed on resource-constrained devices due to their extensive parameters and reliance on dense multiplications, resulting in high memory demands and latency bottlenecks. Shift-and-add reparameterization offers a promising solution by replacing costly multiplications with hardware-friendly primitives in both the attention and multi-layer perceptron (MLP) layers of an LLM. However, current reparameterization techniques require training from scratch or full parameter fine-tuning to restore accuracy, which is resource-intensive for LLMs. To address this, we propose accelerating pretrained LLMs through post-training shift-and-add reparameterization, creating efficient multiplication-free models, dubbed ShiftAddLLM. Specifically, we quantize each weight matrix into binary matrices paired with group-wise scaling factors. The associated multiplications are reparameterized into (1) shifts between activations and scaling factors and (2) queries and adds according to the binary matrices. To reduce accuracy loss, we present a multi-objective optimization method to minimize both weight and output activation reparameterization errors. Additionally, based on varying sensitivity across layers to reparameterization, we develop an automated bit allocation strategy to further reduce memory usage and latency. Experiments on five LLM families and eight tasks consistently validate the effectiveness of ShiftAddLLM, achieving average perplexity improvements of 5.6 and 22.7 points at comparable or lower latency compared to the most competitive quantized LLMs at 3 and 2 bits, respectively, and more than 80% memory and energy reductions over the original LLMs. Codes and models are available at https://github.com/GATECH-EIC/ShiftAddLLM.
@@ 1 Introduction
We identify a promising yet unexplored opportunity for improving LLM efficiency: reparameterizing their extensive multiplications with more cost-effective hardware substitutes, such as bitwise shifts and adds. Inspired by practices in computer architecture and digital signal processing, replacing multiplications with bitwise shifts and adds [60, 21] can offer up to 31 energy and 26 area reductions (see Tab. 1). This hardware-inspired approach can lead to efficient and fast implementations, as shown by previous research on ShiftAddNet [63, 64, 66]. Unlike previous reparameterization techniques that require training from scratch or extensive fine-tuning, we propose a new method to integrate the shift-and-add concept into LLMs through post-training optimization.
@@
To design multiplication-less LLMs, we need to address three key challenges: First, how can we effectively reparameterize pretrained LLMs with shifts and adds in a post-training manner? Previous reparameterization techniques [63, 66] often result in significant quantization errors, requiring extensive fine-tuning or retraining to avoid accuracy drops. We aim to develop a ready-to-use post-training reparameterization method for LLMs. Second, how can we mitigate the accuracy drop from shift-and-add reparameterization? Approximating original multiplications with lower-bit shifts and adds typically reduces accuracy. Most studies resort to fine-tuning or increasing model sizes, complicating LLM deployment. We hypothesize that optimizing both weight and activation errors can minimize overall reparameterization error, aligning with recent activation-aware weight quantization methods in LLMs. Third, how can we handle varying sensitivities to reparameterization across different layers and blocks in LLMs? An automated strategy to determine the optimal number of bits for reparameterized weights in each layer is needed. More vulnerable layers should have higher-bit representations, while less sensitive layers can use lower-bit representations. This ensures no bottlenecked layers due to aggressive reparameterization and maximizes redundancy exploitation. To the best of our knowledge, this is the first attempt to address these three challenges for multiplication-less LLMs through post-training reparameterization. Our contributions are summarized as follows:
@@
Our extensive results across five LLMs and eight tasks consistently show the superior accuracy and efficiency trade-offs achieved by ShiftAddLLM, with average perplexity improvements of 5.6 and 22.7 at comparable or even lower latency compared to the most competitive quantized LLMs at three and two bits, respectively, and more than 80% memory and energy reductions over the original LLMs.
@@ 3 Preliminaries
Shift and Add Primitives. Direct hardware implementation of multiplications is often inefficient. Using shift and add operations as “shortcuts” provides a more efficient alternative. Shifts, which are equivalent to multiplying by powers of two, offer a non-uniform quantization solution and can result in significant savings. For example, reductions of 23.8 and 22.3 in energy and area, respectively, as shown in Tab. 1. Similarly, additions can achieve up to 31.0 and 25.5 savings in energy and area. Both primitives have inspired many innovations in efficient model innovations [7, 15, 63, 66].
@@ 4.1 ShiftAddLLM: Post-training Reparameterization of LLMs with Shift and Add Primitives
After the above reparameterization, we can then replace the associated multiplication between weights and activations into two steps: (1) Bitwise shifts between activations and scaling factors. Note that the activation is still in the FP16 format, and the multiplication between a floating-point number and a positive or negative PoT integer can be efficiently implemented by an integer addition instruction on existing hardware following DenseShift [32], as also illustrated in Fig. 1 (c); (2) Queries and adds intermediate shifted activations with the binary matrices. To implement this efficiently and reduce redundant additions or accumulations, as shown in Fig. 1 (d), we pre-compute 256 () possible values for every eight elements in the shifted activations to construct LUTs. This allows us to handle the multiplication between the binary matrix and the shifted activations as queries to the LUTs. We then add all the partial sums to obtain the final output activations. Such LUTs are well supported by existing GPU kernels [44, 27]. The reparameterization can be applied to all weights in pretrained LLMs in a post-training manner, replacing costly multiplications with efficient hardware operations.
Insights. ShiftAddLLM presents a novel multiplication-less approach that leverages non-uniform quantization via BCQ and additive PoT. This methodology enhances the representation capacity for outlier weights and activations of large magnitude compared to uniform quantization. Moreover, additive PoT effectively resolves the issue of limited quantization resolution for non-outlier weights and activations. Overall, it allows the quantization levels to better align with the data distribution.
@@ 4.2 ShiftAddLLM: Multi-objective Optimization
Accuracy vs. Latency Tradeoffs. The column-wise scaling factor design significantly boosts accuracy after reparameterization. However, it does not fully leverage BCQ [44, 27], which process eight elements per row of weights in parallel as LUT keys, resulting in latency overhead for models with 30B parameters. For example, testing on the OPT-30B [67] model and WikiText-2 dataset [37] showed 6.7 perplexity reduction but with a 24.7% latency overhead, as shown in Fig. 3 (b).
@@
Insights. Our multi-objective optimization approach integrates both weight and activation objectives, reducing weight quantization error in an activation-aware manner and output activation error reduction in a weight-aware manner. This synergy, achieved through a simple column-wise or block-wise design, significantly boosts the accuracy of weight-only quantization. This aligns with the principles of previous activation-aware weight quantization methods [34].
@@ 4.3 ShiftAddLLM: Mixed and Automated Bit Allocation
Sensitivity Analysis. We analyze the sensitivity of different layers and blocks in LLMs to shift-and-add reparameterization. As shown in Fig. 4, later blocks incur more quantization or reparameterization errors. Within each block, Q/K linear layers are generally more sensitive to reparameterization than other linear layers. This diverse sensitivity motivates us to explore mixed bit allocations for LLM reparameterization and develop strategies to automatically determine the optimal bit allocations given the average bit budgets.
@@
Insights. Using mixed bits instead of static ones can improve the accuracy-efficiency tradeoffs by adapting the varying sensitivities across layers, e.g., Q/K linear layers exhibit higher sensitivity to reparameterization; Our adopted criteria provide a quick estimation of the reparameterization error.
@@ 5.1 Experiment Settings
Models. We consider five representative SOTA LLM families, including OPT [67], LLaMA-1/2/3 [52, 2], Gemma [38], Mistral [29], and Bloom [45]. Tasks and Datasets. We evaluate all five LLMs on the commonly adopted language modeling task using the WikiText-2 [37] dataset for perplexity measurement. Additionally, we extend the evaluation of the two largest models, OPT-66B and LLaMA-2-70B, to seven downstream tasks for zero-shot task accuracy evaluation. These tasks include ARC (Challenge/Easy) [4], BoolQ [9], Copa [1], PIQA [50], RTE [11], and StoryCloze [39]. Baselines. We consider four SOTA LLM quantization methods: OPTQ [17], LUT-GEMM [44], QuIP [6], and AWQ [34]. Evaluation Metrics. We evaluate ShiftAddLLM and the baselines using both accuracy and efficiency metrics. For accuracy, we evaluate perplexity on the WikiText-2 dataset and zero-shot accuracy on seven downstream tasks. For efficiency, we measure the latency on a single A100-80GB GPU (PCIe) [41] and estimate the energy costs using an Eyeriss-like hardware accelerator [8, 68], which calculates not only computational but also data movement energy.
@@ 5.2 ShiftAddLLM over SOTA LLM Quantization Baselines
Results on OPT Models. To evaluate the effectiveness of our ShiftAddLLM, we compare with four SOTA LLM quantization baselines: OPTQ [17], LUT-GEMM [44], QuIP [6], and AWQ [34]. Using the OPT model family [67] and the WikiText-2 dataset [37], we assess perplexity, GPU latency, and energy costs. As shown in Tab. 2, Ours (Acc.) consistently outperforms all baselines, achieving an average perplexity reduction of 5.63/38.47/5136.13 compared to OPTQ, LUT-GEMM, and AWQ, respectively, at 3 bits. At 2 bits, where most baselines fail with significantly high perplexity, our method maintains low perplexity, and achieves an average 22.74 perplexity reduction over the most competitive QuIP. Also, as shown in Fig. 6 (a & b), Ours (Lat.) consistently achieves better accuracy-latency tradeoffs, achieving 0.91103830.45 perplexity reduction at comparable latency or 6.5%60.1% latency reductions and 26.0%44.7% energy savings at similar or even lower perplexity. Complete quantitative data on accuracy, latency, and energy are provided in Appendix A.
@@
Results on Gemma/Mistral/Bloom Models. We also evaluate ShiftAddLLM on Gemma [38], Mistral [29], and Bloom [45] models, which are among the most popular open-source LLMs and Mixture-of-Expert (MoE) models. As shown in Tab. 4, Ours (Acc.) achieves perplexity reductions of 11.12/29.4 for Gemma-2B, 1.67/16.76 for Mistral-7B, and 3.30/6.93 and 1.76/5.58 for BLOOM-3B/7B, respectively, compared to OPTQ and LUT-GEMM. As shown in Fig. 6 (e), Ours (Lat.) shows better accuracy-latency tradeoffs, e.g., achieving 9.56 perplexity reduction and 11% latency reductions over the OTPQ baseline on Gemma models. Tehse results on five LLM families consistently validate the effectiveness of our ShiftAddLLM.
| Models | Methods | Bits | ARC_C | ARC_E | Copa | BoolQ | PIQA | Storycloze | RTE | Mean |
| OPT-66B | OPTQ [17] | 3 | 24.66 | 48.86 | 70 | 52.05 | 64.47 | 67.09 | 53.07 | 51.99 |
| LUT-GEMM [44] | 3 | 24.15 | 51.85 | 81 | 53.52 | 61.97 | 60.60 | 48.74 | 51.97 | |
| Ours (Acc.) | 3 | 35.24 | 70.88 | 87 | 72.45 | 77.64 | 77.15 | 63.18 | 65.58 | |
| LLaMA-2-70B | OPTQ [17] | 3 | 45.82 | 76.34 | 90 | 81.74 | 79.71 | 77.34 | 67.51 | 74.07 |
| LUT-GEMM [44] | 3 | 47.70 | 76.42 | 89 | 80.31 | 80.20 | 77.78 | 68.59 | 74.29 | |
| Ours (Acc.) | 3 | 48.38 | 77.06 | 93 | 84.25 | 80.47 | 78.49 | 75.09 | 76.68 | |
Zero-shot Downstream Tasks. We extend our evaluation to zero-shot downstream datasets for a more comprehensive assessment. As shown in Tab. 5, Ours (Acc.) consistently improves performance over previous SOTA baselines, achieving an average accuracy gain of 13.59/13.61 and 2.61/2.39 over OPTQ and LUT-GEMM baselines at 3 bits when evaluated on OPT-66B and LLaMA-2-70B, respectively. These experiments demonstrate that our method not only reduce perplexity but also improves downstream task accuracy.

### v4

@@ Abstract
Large language models (LLMs) have shown impressive performance on language tasks but face challenges when deployed on resource-constrained devices due to their extensive parameters and reliance on dense multiplications, resulting in high memory demands and latency bottlenecks. Shift-and-add reparameterization offers a promising solution by replacing costly multiplications with hardware-friendly primitives in both the attention and multi-layer perceptron (MLP) layers of an LLM. However, current reparameterization techniques require training from scratch or full parameter fine-tuning to restore accuracy, which is resource-intensive for LLMs. To address this, we propose accelerating pretrained LLMs through post-training shift-and-add reparameterization, creating efficient multiplication-free models, dubbed ShiftAddLLM. Specifically, we quantize each weight matrix into binary matrices paired with group-wise scaling factors. The associated multiplications are reparameterized into (1) shifts between activations and scaling factors and (2) queries and adds according to the binary matrices. To reduce accuracy loss, we present a multi-objective optimization method to minimize both weight and output activation reparameterization errors. Additionally, based on varying sensitivity across layers to reparameterization, we develop an automated bit allocation strategy to further reduce memory usage and latency. Experiments on five LLM families and eight tasks consistently validate the effectiveness of ShiftAddLLM, achieving average perplexity reductions of 5.6 and 22.7 points at comparable or lower latency compared to the most competitive quantized LLMs at 3- and 2-bit precision, respectively, and more than 80% memory and energy reductions over the original LLMs. Codes and models are available at https://github.com/GATECH-EIC/ShiftAddLLM.
@@ 1 Introduction
We identify a promising yet unexplored opportunity for improving LLM efficiency: reparameterizing their extensive multiplications with more cost-effective hardware substitutes, such as bitwise shifts and adds. Inspired by practices in computer architecture and digital signal processing, replacing multiplications with bitwise shifts and adds [66, 22] can offer up to 31 energy and 26 area reductions (see Tab. 1). This hardware-inspired approach can lead to efficient and fast implementations, as shown by previous research on ShiftAddNet [69, 70, 72]. Unlike previous techniques that require training from scratch or extensive fine-tuning, we propose a new method to integrate the shift-and-add concept into LLMs through post-training optimization.
@@
To design multiplication-less LLMs, we need to address three key challenges: First, how can we effectively reparameterize pretrained LLMs with shifts and adds in a post-training manner? Previous reparameterization techniques [69, 72] can result in nontrivial quantization errors, requiring fine-tuning or retraining to avoid accuracy drops. We aim to develop a ready-to-use post-training reparameterization method for LLMs. Second, how can we mitigate the accuracy drop from shift-and-add reparameterization? Approximating original multiplications with lower-bit shifts and adds typically reduces model accuracy. Most studies resort to fine-tuning or increasing model sizes, complicating LLM deployment. We hypothesize that optimizing both weight and activation errors can minimize overall reparameterization error, aligning with recent activation-aware weight quantization methods in LLMs. Third, how can we handle varying sensitivities to reparameterization across different layers and blocks in LLMs? An automated strategy to determine the optimal number of bits for reparameterized weights in each layer is needed. More vulnerable layers should have higher-bit representations, while less sensitive layers can use lower-bit representations. This ensures no bottlenecked layers due to aggressive reparameterization and maximizes redundancy exploitation. To the best of our knowledge, this is the first attempt to address these three challenges for multiplication-less LLMs through post-training reparameterization. Our contributions are summarized as follows:
@@
Our extensive results across five LLMs and eight tasks consistently show the superior accuracy and efficiency trade-offs achieved by ShiftAddLLM, with average perplexity reductions of 5.6 and 22.7 at comparable or even lower latency compared to the most competitive quantized LLMs at three and two bits, respectively, and more than 80% memory and energy reductions over the original LLMs.
@@ 3 Preliminaries
Shift and Add Primitives. Direct hardware implementation of multiplications is often inefficient. Using shift and add operations as “shortcuts” provides a more efficient alternative. Shifts, which are equivalent to multiplying by powers of two, offer a non-uniform quantization solution and can result in significant savings. For example, we tested matrix multiplication from one MLP layer of OPT-66B between weight and activation using FP16 MACs and our 3-bit ShiftAddLLM. Energy consumption was 80.36J vs. 9.77J, achieving 87.8% savings with our method. Both primitives have inspired many innovations in efficient model innovations [7, 16, 69, 72].
@@ 4.1 ShiftAddLLM: Post-training Reparameterization of LLMs with Shift and Add Primitives
After the above reparameterization, we can then replace the associated multiplication between weights and activations into two steps: (1) Bitwise shifts between activations and scaling factors. Note that the activation is still in the FP16 format, and the multiplication between a floating-point number and a positive or negative PoT integer can be efficiently implemented by an integer addition instruction on existing hardware following DenseShift [36], as also illustrated in Fig. 1 (c); (2) Queries and adds intermediate shifted activations with the binary matrices. To implement this efficiently and reduce redundant additions or accumulations, as shown in Fig. 1 (d), we pre-compute 256 () possible values for every eight elements in the shifted activations to construct LUTs. Here every eight grouped binary weights form an 8-bit key. Suppose the shifted activation is an -dimensional vector. In that case, we will get LUTs, where the grouped binary weights are used as keys, and the precomputed partial sums are stored as values. This allows us to handle the multiplication between the binary matrix and the shifted activations as queries to the LUTs. We then add all the partial sums to obtain the final output activations in FP16 format. Such LUTs are well supported by existing GPU kernels [48, 29]. The reparameterization can be applied to all weights in pretrained LLMs in a post-training manner, replacing costly multiplications with efficient hardware operations.
Takeaway. ShiftAddLLM presents a novel multiplication-less approach that leverages non-uniform quantization via BCQ and additive PoT. This methodology enhances the representation capacity for outlier weights and activations of large magnitude compared to uniform quantization. Moreover, additive PoT effectively resolves the issue of limited quantization resolution for non-outlier weights and activations. Overall, it allows the quantization levels to better align with the data distribution.
@@ 4.2 ShiftAddLLM: Multi-objective Optimization
Accuracy vs. Latency Tradeoffs. The column-wise scaling factor design significantly boosts accuracy after reparameterization. However, it does not fully leverage BCQ [48, 29], which process eight elements per row of weights in parallel as LUT keys, resulting in latency overhead for models with 30B parameters. For example, testing on the OPT-30B [74] model and WikiText-2 dataset [41] showed (16.3 9.6) 6.7 perplexity reduction but with a 24.7% latency overhead, as shown in Fig. 3 (b).
@@
Takeaway. Our multi-objective optimization approach integrates both weight and activation objectives, reducing weight quantization error in an activation-aware manner and output activation error reduction in a weight-aware manner. This synergy, achieved through a simple column-wise or block-wise design, significantly boosts the accuracy of weight-only quantization. This aligns with the principles of previous activation-aware weight quantization methods [38].
@@ 4.3 ShiftAddLLM: Mixed and Automated Bit Allocation
Sensitivity Analysis. We analyze the sensitivity of different layers and blocks in LLMs to shift-and-add reparameterization. As shown in Fig. 4, later blocks incur more quantization or reparameterization errors. Within each block, Query/Key (Q/K) layers are generally more sensitive to reparameterization than other linear layers. This diverse sensitivity motivates us to explore mixed bit allocations for LLM reparameterization and develop strategies to automatically determine the optimal bit allocations given the average bit budgets.
@@
Takeaway. Using mixed bits instead of static ones can improve the accuracy-efficiency tradeoffs by adapting the varying sensitivities across layers, e.g., Q/K linear layers exhibit higher sensitivity to reparameterization; Our adopted criteria provide a quick estimation of the reparameterization error.
@@ 5.1 Experiment Settings
Models. We consider five representative SOTA LLM families, including OPT [74], LLaMA-1/2/3 [58, 2], Gemma [42], Mistral [31], and Bloom [49]. Tasks and Datasets. We evaluate all five LLMs on the commonly adopted language modeling task using the WikiText-2 [41] dataset for perplexity measurement. Additionally, we extend the evaluation of the two largest models, OPT-66B and LLaMA-2-70B, to eight downstream tasks for zero-shot accuracy evaluation. These tasks include ARC (Challenge/Easy) [4], BoolQ [9], Copa [1], PIQA [56], RTE [11], StoryCloze [43], and MMLU [26]. Baselines. We consider four SOTA LLM quantization methods: OPTQ [18], LUT-GEMM [48], QuIP [6], and AWQ [38]. Evaluation Metrics. We evaluate ShiftAddLLM and the baselines using both accuracy and efficiency metrics. For accuracy, we evaluate perplexity on the WikiText-2 dataset and zero-shot accuracy on eight downstream tasks. For efficiency, we measure the latency on a single A100-80GB GPU (PCIe) [45] and estimate the energy costs using an Eyeriss-like hardware accelerator [8, 75], which calculates not only computational but also data movement energy (within 18% of the differences with Eyeriss’s chip measurement results as claimed).
@@ 5.2 ShiftAddLLM over SOTA LLM Quantization Baselines
Results on OPT Models. To evaluate the effectiveness of our ShiftAddLLM, we compare against four SOTA LLM quantization baselines: OPTQ [18], LUT-GEMM [48], QuIP [6], and AWQ [38]. Using the OPT model family [74] and the WikiText-2 dataset [41], we assess perplexity, GPU latency, and energy costs. As shown in Tab. 2, Ours (Acc.) consistently outperforms all baselines, achieving an average perplexity reduction of 5.63/38.47/5136.13 compared to OPTQ, LUT-GEMM, and AWQ, respectively, at 3 bits. At 2 bits, where most baselines fail with significantly high perplexity, our method maintains low perplexity, and achieves an average 22.74 perplexity reduction over the most competitive QuIP. Also, as shown in Fig. 6 (a & b), Ours (Lat.) consistently achieves better accuracy-latency tradeoffs, with a perplexity reduction of 0.91103830.45 at comparable latency or 6.5%60.1% latency reductions and 26.0%44.7% energy savings at similar or even lower perplexity. Complete quantitative data on accuracy, latency, and energy is provided in Appendix A.
@@
Results on Gemma/Mistral/Bloom Models. We also evaluate ShiftAddLLM on Gemma [42], Mistral [31], and Bloom [49] models, which are among the most popular open-source LLMs and Mixture-of-Expert (MoE) models. As shown in Tab. 4, Ours (Acc.) achieves perplexity reductions of 11.12/29.4 for Gemma-2B, 1.67/16.76 for Mistral-7B, and 3.30/6.93 and 1.76/5.58 for BLOOM-3B/7B, respectively, compared to OPTQ and LUT-GEMM. As shown in Fig. 6 (e), Ours (Lat.) shows better accuracy-latency tradeoffs, e.g., achieving 9.56 perplexity reduction and 11% latency reductions over the OTPQ baseline on Gemma models. These results on five LLM families consistently validate the effectiveness of our ShiftAddLLM.
Zero-shot Downstream Tasks. We extend our evaluation to zero-shot downstream datasets for a more comprehensive assessment. As shown in Tab. 5, Ours (Acc.) consistently improves performance over previous SOTA baselines, achieving an average accuracy gain of 13.37/13.19 and 2.55/2.39 over OPTQ and LUT-GEMM baselines at 3 bits when evaluated on OPT-66B and LLaMA-2-70B, respectively. These experiments demonstrate that our method not only reduces perplexity but also improves downstream task accuracy.
Models | Methods | Bits | ARC_C | ARC_E | Copa | BoolQ | PIQA | Storycloze | RTE | MMLU | Mean |
OPT-66B | Floating Point | 16 | 37.20 | 71.25 | 86 | 69.82 | 78.67 | 77.47 | 60.65 | 25.890.37 | 63.37 |
OPTQ [18] | 3 | 24.66 | 48.86 | 70 | 52.05 | 64.47 | 67.09 | 53.07 | 23.980.36 | 50.52 | |
LUT-GEMM [48] | 3 | 24.15 | 51.85 | 81 | 53.52 | 61.97 | 60.60 | 48.74 | 23.730.36 | 50.70 | |
Ours (Acc.) | 3 | 35.24 | 70.88 | 87 | 72.45 | 77.64 | 77.15 | 63.18 | 27.560.38 | 63.89 | |
LLaMA-2-70B | Floating Point | 16 | 49.57 | 76.14 | 90 | 82.57 | 80.79 | 78.61 | 68.23 | 65.240.37 | 72.89 |
OPTQ [18] | 3 | 45.82 | 76.34 | 90 | 81.74 | 79.71 | 77.34 | 67.51 | 60.140.36 | 72.33 | |
LUT-GEMM [48] | 3 | 47.70 | 76.42 | 89 | 80.31 | 80.20 | 77.78 | 68.59 | - | - | |
Ours (Acc.) | 3 | 48.38 | 77.06 | 93 | 84.25 | 80.47 | 78.49 | 75.09 | 62.330.38 | 74.88 | |
@@
Results of Mixed Bit Allocation. We evaluate our mixed bit allocation strategy (see Sec. 4.3) and compare Ours (Mixed) with Ours (Lat.). As shown in Tab. 6, Ours (Mixed) further reduces the perplexity by an average of 79.45 for OPT model families under comparable or even less latency. We provide more results in Appendix F to validate the effectiveness of our mixed bit allocation strategy.